OpenStudy (anonymous):

PLS HELP ALGEBRA 1 AND I DONT EVEN KNOW WHERE TO START

OpenStudy (anonymous):

STEP 1. Choose one of the groups of equations from the table below. Indicate which group you choose in your word processing file. Group 1 y = negative three-halvesx + 4 –3x + 2y = 8 Group 2 y = negative 2 over 3x – 2 –2x + 3y = -6 Group 3 y = negative 1 over 3x + 3 –x + 3y = 9 Group 4 y = negative 1 over 6x + 5 –x + 6y = 30 STEP 2. Type the first equation from the group you selected and identify the slope and the y-intercept of the equation. STEP 3. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. STEP4. Type the second equation from the group you selected and change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation. STEP5 .Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. STEP 6. In a paragraph of at least three sentences, explain how the graphs of the two equations in your group are the same and how they are different. Be sure to use complete sentences.

OpenStudy (anonymous):

i want to work with group 3

OpenStudy (espex):

Then to start you take your equation, y = negative 1 over 3x + 3 –x + 3y = 9, and put it into slope-intercept form.

OpenStudy (espex):

To identify your y intercept you simply plug 0 in for your x. Given a starting point of (0,y) you can use your slope, which is rise over run, you can describe how many steps over and up you will need for your next point.

OpenStudy (espex):

Is it safe to presume that your two equations for group 3 are \[y=-\frac{1}{3}x+3\] and \[-x+3y=9\]

OpenStudy (espex):

Put each of those equations into slope intercept form (y=mx+b) and solve for your y intercept to complete step 2.